# Title of Your Paper

### v1

@@
[…]
Title of Your Paper
Hana Samad Responsible AI Lab, National Fair Housing Alliance, Washington, DC, United States Michael Akinwumi Corresponding author: makinwumi@nationalfairhousing.org Responsible AI Lab, National Fair Housing Alliance, Washington, DC, United States Jameel Khan Housing & Community Development, National Fair Housing Alliance, Washington, DC, United States Christoph Mügge-Durum Responsible AI Lab, National Fair Housing Alliance, Washington, DC, United States Emmanuel O. Ogundimu Department of Mathematical Sciences, Durham University, Durham, United Kingdom
Selecting for Less Discriminatory Algorithms: A Relational Search Framework for Navigating Fairness-Accuracy Trade-offs in Practice
Hana Samad Responsible AI Lab, National Fair Housing Alliance, Washington, DC, United States Michael Akinwumi Corresponding author: makinwumi@nationalfairhousing.org Responsible AI Lab, National Fair Housing Alliance, Washington, DC, United States Jameel Khan Housing & Community Development, National Fair Housing Alliance, Washington, DC, United States Christoph Mügge-Durum Responsible AI Lab, National Fair Housing Alliance, Washington, DC, United States Emmanuel O. Ogundimu Department of Mathematical Sciences, Durham University, Durham, United Kingdom
@@
Against these considerations, the paper revisits Lee and Floridi (2021)’s relational fairness approach using updated 2021 Home Mortgage Disclosure Act (HMDA) data, and proposes an expansion of the scope of the LDA search process. We argue that extending the LDA search horizontally, considering fairness across model families themselves, provides a lightweight complement, or alternative, to within-model hyperparameter optimization, when operationalizing fairness in non-experimental, resource constrained settings. Fairness metrics alone offer useful, but insufficient signals to accurately evaluate candidate LDAs. Rather, by using a horizontal LDA search approach with a relational trade-off framework, we demonstrate a responsible minimum viable LDA search on real-world lending outcomes. Organizations can modify this approach to systematically compare, evaluate, and select LDAs that optimize fairness and accuracy in a sector-based contextualized manner.
1 Introduction
@@
Discriminatory policies and exclusionary laws fundamentally skewed generational access to financial credit and housing opportunities, thus people of color are inherently underrepresented in financial service and homeownership data. 2021 Home Mortgage Disclosure Act (HMDA) data reflects this idea, revealing significant racial disparities in homeownership rates. The rates found in these foundational datasets are the basis for algorithmic system training, informing current access and data-driven material allocation decisions. Unlike humans, who can demonstrate empathy and perhaps adjust decisions accordingly (see Das et al. (2023)), algorithms operate with less transparency and are not inherently equipped to modify or contextualize their outputs in the face of ethical considerations or sociohistorical context. We refer interested readers to learn more about this algorithmic dilemma in relation to the nuances of the housing context, as well as the associated history of discrimination in the following articles ( (United States, 1988),(Federal Financial Institutions Examination Council, 2023), (Lee, 2018), (Urban Institute, 2020), (Draude et al., 2020), and Section 2 of (Lee and Floridi, 2021)).
Although this study focuses primarily on mortgage lending and housing, the issues and questions explored are directly applicable to other sectors grappling with the same journeys and impediments to adopting automation. Choosing to offer AI-based services and products has serious implications for the risk management processes of private sector entities throughout the product lifecycle, including model procurement, development, and compliance to mitigate consumer harm. What liability exists in AI-based products and where opportunities for risk management exist in the development pipeline are the key questions on industry’s mind. This in practice—a lack of actionable harm mitigation strategies, frameworks, and common industry standards—directly results in industry hesitance to implement algorithmic systems within federally flagged high-impact areas (Office of Management and Budget, 2025, p. 21).
@@
In this work, we focus on an inference stage, trade-off curve analysis technique that Lee and Floridi (2021) developed and applied to 2011 HMDA data. The methodology was framed as a balance of objectives rather than an absolute mathematical fairness criterion; we apply this technique to more recently available 2021 HMDA data, ten years later. This longitudinal approach aims to assess the robustness of the relational fairness tradeoff approach across current patterns in mortgage data in America, testing if the methodology: 1) can address the contextual complexity of bias and fairness in this area, and 2) function as a clear, low barrier framework that public and private decision-makers can employ to select fairer algorithmic systems.
This study adds new insights on three key issues by introducing the framework to ongoing legal, ML, and algorithmic governance discussions, namely the utility of model multiplicity and LDAs in practice. First, we contribute to the growing LDA literature, particularly existing work on within-model hyperparameter optimization approaches ((Gillis et al., 2024); (Laufer et al., 2025))—what we refer to as vertical LDA searches—by arguing that the scope of the LDA search itself should be expanded to include testing model families themselves initially as a key fairness consideration. This process we refer to as a horizontal LDA search—an early-stage model development intervention that exploits structure-level algorithmic differences for improved sector-dependent fairness gains. Institutions can reproduce this internally within their existing development processes to achieve fairness and compliance in a lightweight, scalable, and reasonable manner. Secondly, we explore the relational approach in navigating the intersection of financial inclusion and racial equity within mortgage lending practices. We highlight the effects of including or excluding race as a feature in algorithmic decision-making in lending, evaluating how this impacts real-world material outcomes and predictive performance, in a sector where the foundational data is fundamentally compromised. Finally, we take these ideas in tandem and translate them into recommendations that industry and regulatory actors can focus on to successfully encourage the broader adoption of LDAs as a paradigm for legible, compliance-driven, and socially responsible AI development.
@@ 2 Literature Review
In trying to understand and evaluate these separate methodological tensions, this literature review can be broken down into the following parts: 1) Understanding the conceptual landscape of fairness in machine learning, 2) Discussing fairness metrics, their use cases for selecting fairer models, 3) Evaluating the relational approach to fairness-accuracy trade-offs as the basis for a potential framework in real-world cases, 4) Reviewing common machine learning algorithms, and 5) Examining the applicability of model multiplicity and scope of LDAs.
@@
2.1 Defining Fairness for Algorithmic Decision-making
@@
The increasing deployment of algorithmic decision-making systems across various sectors has brought concerns about fairness and potential discrimination to the forefront. Particularly in sensitive areas like mortgage lending, there is a risk that algorithms, often trained on historically biased data, may perpetuate or even amplify existing societal inequalities (Leininger et al., 2025). In response, the field of algorithmic fairness has burgeoned, proposing numerous mathematical formalizations intended to capture different notions of fairness. However, many of these definitions are presented as absolute, universally applicable conditions, potentially overlooking the intricate realities of specific applications.
@@
Lee and Floridi (2021) challenge this paradigm, arguing that the practical and ethical trade-offs inherent in complex decision-making processes are unavoidable. They propose a shift towards understanding fairness not as a static, absolute mathematical state to be achieved, but as a relational notion. This perspective advocates for evaluating the fairness of an algorithmic system relative to other viable options and explicitly considering the trade-offs between various, often competing, objectives. Applying this to 2011 Home Mortgage Disclosure Act (HMDA) data, they emphasize the limitations of standard mathematical metrics in capturing the nuanced ethical dilemmas faced by lenders at face value. This is particularly important for lending given biases embedded within legacy data, and the difficulty of untangling legitimate risk proxies from potentially discriminatory proxies for protected characteristics, e.g., race. The relational approach encourages an evaluation focused on achieving a model as close to equilibrium between desired outcomes, such as enhancing financial access, and potential negative impacts on specific groups. It marks a significant conceptual move, steering the literature away from the pursuit of a singular, perfect mathematical fairness definition towards a more pragmatic, context-dependent framework centered on navigating and managing competing values.
@@
Algorithmic fairness metrics are broadly categorized under two main paradigms: individual fairness and group fairness (Anderson and Visweswaran, 2025). Individual fairness (IF) adheres to the principle that ”similar individuals should be treated similarly”. In this context, similarity is determined based on task-relevant characteristics, independent of protected attributes. For instance, two loan applicants with comparable financial profiles should ideally receive similar loan decisions. However, defining an appropriate, context-specific similarity metric remains a significant challenge, and subjective or data-driven definitions risk encoding existing biases. Group fairness (GF), conversely, focuses on achieving parity in outcomes or statistical measures across different demographic groups, typically defined by protected attributes such as race, sex, religion, national origin, or familial status (Gohar and Cheng, 2023). The objective is often to equalize specific rates, like approval rates or error rates, between these groups (Zhou, 2022).
@@
A critical tension exists between these two paradigms: satisfying group fairness criteria does not guarantee individual fairness, and enforcing group-level parity can sometimes necessitate treating similar individuals differently based on their group membership, directly conflicting with the IF principle (Zhou, 2022). This conflict often reflects a deeper ethical divergence between ensuring procedural consistency for individuals (treating like cases alike) and pursuing distributive justice to rectify group-level disparities, which may be rooted in historical bias reflected in the data. Furthermore, focusing solely on isolated protected attributes can obscure intersectional discrimination experienced by individuals with multiple identities (e.g., race and gender) (Gohar and Cheng, 2023).
@@
To address identified biases and strive towards fairness goals, various intervention techniques have been developed, typically categorized according to the stage of the machine learning pipeline at which they are applied: pre-processing, in-processing, and post-processing (Dieng et al., 2025).
  1. (a)
@@
    Pre-processing techniques: These techniques involve modifying the training data before a model is trained, removing or mitigating biases present in the data itself. Examples include suppressing sensitive attributes or their proxies, re-labeling historically biased outcomes, re-weighting data points to balance group representation, or learning on transformed data representations designed to be fair Nielsen (2020). The main advantage is that these techniques are model-agnostic and tackle bias at its source; some research suggests that successful pre-processing to approximate an unbiased world could resolve fairness-accuracy trade-offs (Leininger et al., 2025). However, pre-processing can significantly alter the original data, simple suppression is often ineffective due to correlations, and relabeling can be subjective and practically challenging.
  2. (b)
@@
    In-processing techniques: They integrate fairness considerations directly into the model training process, typically by modifying the learning algorithm or its objective function (Leininger et al., 2025). Common approaches include adding fairness-related regularization terms to the loss function to penalize unfair predictions (Petersen et al., 2021) or imposing fairness constraints during optimization. In-processing methods can potentially achieve a better balance between fairness and accuracy by optimizing them jointly (Zemel et al., 2024). These in- processing methods include in-training approaches such as adversarial debiasing (Zhang et al., 2018) and distribution matching (Quadrianto and Sharmanska, 2017) techniques, the latter of which we previously applied to mortgage data. [^1] We applied the distribution matching technique to mortgage underwriting and pricing models and it resulted in reduced racial disparities in loan approvals by 13% and decreased borrowing costs for Black and Hispanic applicants by 20% without compromising system accuracy. These methods promote fairness while maintaining efficiency. See https://nationalfairhousing.org/wp-content/uploads/2024/04/Unlocking-Fairness-Final_April-2024.pdf Drawbacks include often being model-specific, requiring access to and modification of the training pipeline, and incurring the computational cost of retraining (Zemel et al., 2024).
  3. (c)
@@
    Post-processing techniques: They adjust the outputs of an already trained model without altering the model itself (Zemel et al., 2024). Examples include setting different classification thresholds for different groups or learning a separate transformation function for the model’s scores. The primary appeal lies in being model-agnostic, avoiding costly retraining, and applicability to deployed or black-box systems. However, their effectiveness is constrained by the information contained in the original model’s predictions, they may be less impactful than in-processing methods, and can sometimes improve group fairness at the expense of individual fairness or accuracy.
@@
The choice among these intervention strategies is not merely technical but involves navigating practical constraints—such as access to data, control over the model training process, and computational resources—alongside strategic goals regarding the desired fairness definition and tolerance for accuracy trade-offs (Leininger et al., 2025).
2.2 Quantifying Fairness: A Review of Mathematical Fairness Metrics
@@
In conjunction with theoretical fairness discussions, the field has created a toolkit of fairness metrics to quantify and describe variability between different subgroups within a dataset—-in the housing sector this usually includes demographic parity, standardized mean deviation, disparate/adverse impact, and mean absolute deviation. A common definition for fairness in this context points to opportunity access rather than to outcomes themselves Lee and Floridi (2021). Equalization of evaluation metrics attempt to formalize this belief that a model’s predictive accuracy should not vary between racial groups. The most utilized and cited of these include equal opportunity (EOP), false positive error rate balance, equal odds, positive predictive parity, positive class balance, and negative class balance.
@@
Equal opportunity, as discussed in Hardt et al. (2016) ensures equal chances of loan or housing approval for all credit-worthy applicants, focusing on minimizing false negative rates across demographics. Chouldechova (2017)’s work on false positive error rate balance seeks to ensure that similar non-credit worthy applicants are equally likely to be denied loans, preventing bias in loan disapproval. Equal odds requires algorithms to achieve equal true positive and false positive rates across groups, ensuring accuracy in identifying credit-worthy and non-worthy applicants, irrespective of demographic background Hardt et al. (2016). Positive predictive parity, aims for equal probabilities of correct loan repayment predictions across different groups, ensures that the algorithm’s trust is uniformly distributed, preventing skewed good faith predictions that favor specific demographics. Finally, Kleinberg et al. (2016)’s work on positive class balance and negative class balance ensures that the confidence in predictions—whether for loan repayment or default—is fairly balanced across groups, reflecting an objective assessment of an applicant’s financial behavior.
@@
Collectively, fairness metrics describe data variance and facilitate the evaluation of outcome distributions. For developers, policymakers, and stakeholders navigating the concerns associated with algorithmic decision-making, they act as a key benchmarking strategy for documenting how ’fairness’ within an algorithm is distributed across subgroups in a dataset. However, as remarked in Lee and Floridi (2021), while adept at quantifying disparities, it is also essential to note that these traditional metrics solely reflect relative comparisons between groups, similar individuals, or group error rates, without necessarily accounting for the biases embedded within the data inherently. This can result then in different effect distributions if comparing the conclusions suggested by metrics with real-world consequences. This is demonstrated in Noriega-Campero et al. (2020)’s applied case study in algorithmic targeting of social policy in Columbia and Costa Rica. While finding overwhelming positive predictive improvements by using AI-based systems compared to the status quo, their analysis underscores that the degree of these improvements were not evenly dispersed. Error reduction ranged between 5.74% to 22.57% in some cases, across various urban-rural, geographic, gender, and familial divisions, highlighting the importance of documenting and testing fairness outcomes across segmented subgroups in high-impact use cases. Fairness metrics, though helpful in highlighting disparities, often present challenges in implementation, particularly in cases when the proxies for outcomes are closely intertwined with those for race, meaning they are best used when heavily contextualized to the data’s use case.
@@
2.3 Critical Evaluation and Adaptation of Lee and Floridi (2021): Strengths and Limitations of the Relational Trade-off Methodology
@@
While informed by the work of Lee and Floridi (2021), the research presented herein adapts certain assumptions employed in their analysis, based on limitations identified within their methodology and the specific requirements of this study. The methodology proposed in Section 5 of the paper offers a distinct approach centered on relational trade-offs. It involves several key steps: first, operationalizing relevant real-world objectives (e.g., defining and quantifying financial inclusion and negative impact on minority borrowers); second, building various predictive algorithms and computing metrics related to these objectives; third, identifying potential proxies for protected characteristics within the model features; and finally, selecting an algorithm by visualizing and evaluating the trade-offs between the competing objectives.
@@
This approach possesses notable strengths. It promotes a holistic and contextual view of fairness, moving beyond abstract mathematical conditions to evaluate algorithms within their specific application context and considering multiple, potentially conflicting, objectives. A significant advantage is its transparency regarding ethical trade-offs; it makes the inherent conflicts between values (like profit versus fairness) explicit and quantifiable, compelling decision-makers to confront and prioritize these trade-offs, aligning with broader calls for tools to evaluate AI trustworthiness (Petersen, 2023). This focus on measurable impacts and comparing concrete algorithmic alternatives lends the methodology practical relevance and actionability for stakeholders, in this case lenders, aiding them in selecting algorithms that align with their ethical commitments and level of risk tolerance. Furthermore, the methodology explicitly incorporates steps to identify and analyze potential proxies for protected attributes, fostering a more profound understanding of how bias might operate, even when sensitive attributes are formally excluded.
@@
The methodology, however, also presents limitations. Its outcomes are highly dependent on the initial assumptions made when operationalizing objectives; how concepts like ”financial inclusion” are defined and measured significantly shapes the resulting analysis. This leads to potential subjectivity, as the interpretation of the trade-offs and the final algorithm choice rely heavily on the decision-maker’s values, potentially lacking the perceived objectivity of purely mathematical metrics. The analysis is also heavily dependent on data quality and availability; Lee and Floridi (2021) themselves noted limitations in the public HMDA data used in their case study, which lacked crucial variables like default outcomes and credit scores, impacting the robustness of the final trade-off evaluation. Moreover, quantifying all relevant factors (e.g., regulatory compliance, long-term societal effects, model interpretability) can be challenging, and the original study explored a limited scope of algorithms and stakeholder perspectives. Ultimately, the framework functions less as a definitive solution to fairness problems and more as a structured process for deliberation. It requires stakeholders to make explicit value judgments at multiple stages—defining objectives, assessing proxies, prioritizing trade-offs—thereby facilitating, rather than automating, ethical decision-making. Its strength lies in clarifying these choices, while its weakness is its reliance on the quality of the inputs and the deliberation itself.

### v2

# Selecting for Less Discriminatory Algorithms: A Relational Search Framework for Navigating Fairness-Accuracy Trade-offs in Practice

@@
Abstract.
@@
Against these considerations, the paper revisits lee21’s relational fairness approach using updated 2021 Home Mortgage Disclosure Act (HMDA) data, and proposes an expansion of the scope of the LDA search process. We argue that extending the LDA search horizontally, considering fairness across model families themselves, provides a lightweight complement, or alternative, to within-model hyperparameter optimization, when operationalizing fairness in non-experimental, resource constrained settings. Fairness metrics alone offer useful, but insufficient signals to accurately evaluate candidate LDAs. Rather, by using a horizontal LDA search approach with a relational trade-off framework, we demonstrate a responsible minimum viable LDA search on real-world lending outcomes. Organizations can modify this approach to systematically compare, evaluate, and select LDAs that optimize fairness and accuracy in a sector-based contextualized manner.
Less Discriminatory Alternatives, LDA search, fairness-accuracy, horizontal LDA search, HMDA, financial services, sector-based, model multiplicity, relational trade-offs, operationalizing fairness
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Computing methodologies Machine learningccs: Applied computing Multi-criterion optimization and decision-makingccs: Applied computing Lawccs: Computing methodologies Search methodologiesccs: Computing methodologies Artificial intelligence
1. Introduction
@@
Discriminatory policies and exclusionary laws fundamentally skewed generational access to financial credit and housing opportunities, thus people of color are inherently underrepresented in financial service and homeownership data. 2021 Home Mortgage Disclosure Act (HMDA) data reflects this idea, revealing significant racial disparities in homeownership rates. The rates found in these foundational datasets are the basis for algorithmic system training, informing current access and data-driven material allocation decisions. Unlike humans, who can demonstrate empathy and perhaps adjust decisions accordingly (see das2023algorithmic), algorithms operate with less transparency and are not inherently equipped to modify or contextualize their outputs in the face of ethical considerations or sociohistorical context. We refer interested readers to learn more about this algorithmic dilemma in relation to the nuances of the housing context, as well as the associated history of discrimination in the following articles ( (us1988),(ffiec_hmda), (TurnerLee2018-TURDRB), (ui2020), (draude2020situated), and Section 2 of (lee21)).
Although this study focuses primarily on mortgage lending and housing, the issues and questions explored are directly applicable to other sectors grappling with the same journeys and impediments to adopting automation. Choosing to offer AI-based services and products has serious implications for the risk management processes of private sector entities throughout the product lifecycle, including model procurement, development, and compliance to mitigate consumer harm. What liability exists in AI-based products and where opportunities for risk management exist in the development pipeline are the key questions on industry’s mind. This in practice—a lack of actionable harm mitigation strategies, frameworks, and common industry standards—directly results in industry hesitance to implement algorithmic systems within federally flagged high-impact areas (OMB2025AI, p. 21).
@@
In this work, we focus on an inference stage, trade-off curve analysis technique that lee21 developed and applied to 2011 HMDA data. The methodology was framed as a balance of objectives rather than an absolute mathematical fairness criterion; we apply this technique to more recently available 2021 HMDA data, ten years later. This longitudinal approach aims to assess the robustness of the relational fairness tradeoff approach across current patterns in mortgage data in America, testing if the methodology: 1) can address the contextual complexity of bias and fairness in this area, and 2) function as a clear, low barrier framework that public and private decision-makers can employ to select fairer algorithmic systems.
This study adds new insights on three key issues by introducing the framework to ongoing legal, ML, and algorithmic governance discussions, namely the utility of model multiplicity and LDAs in practice. First, we contribute to the growing LDA literature, particularly existing work on within-model hyperparameter optimization approaches ((gillis2024); (laufer2025))—what we refer to as vertical LDA searches—by arguing that the scope of the LDA search itself should be expanded to include testing model families themselves initially as a key fairness consideration. This process we refer to as a horizontal LDA search—an early-stage model development intervention that exploits structure-level algorithmic differences for improved sector-dependent fairness gains. Institutions can reproduce this internally within their existing development processes to achieve fairness and compliance in a lightweight, scalable, and reasonable manner. Secondly, we explore the relational approach in navigating the intersection of financial inclusion and racial equity within mortgage lending practices. We highlight the effects of including or excluding race as a feature in algorithmic decision-making in lending, evaluating how this impacts real-world material outcomes and predictive performance, in a sector where the foundational data is fundamentally compromised. Finally, we take these ideas in tandem and translate them into recommendations that industry and regulatory actors can focus on to successfully encourage the broader adoption of LDAs as a paradigm for legible, compliance-driven, and socially responsible AI development.
@@ 2. Literature Review
In trying to understand and evaluate these separate methodological tensions, this literature review can be broken down into the following parts: 1) Understanding the conceptual landscape of fairness in machine learning, 2) Discussing fairness metrics, their use cases for selecting fairer models, 3) Evaluating the relational approach to fairness-accuracy trade-offs as the basis for a potential framework in real-world cases, and 4) Examining the applicability of model multiplicity and scope of LDAs.
@@
2.1. Defining Fairness for Algorithmic Decision-making
@@
The increasing deployment of algorithmic decision-making systems across various sectors has brought concerns about fairness and potential discrimination to the forefront. Particularly in sensitive areas like mortgage lending, there is a risk that algorithms, often trained on historically biased data, may perpetuate or even amplify existing societal inequalities (lein). In response, the field of algorithmic fairness has burgeoned, proposing numerous mathematical formalizations intended to capture different notions of fairness. However, many of these definitions are presented as absolute, universally applicable conditions, potentially overlooking the intricate realities of specific applications.
@@
lee21 challenge this paradigm, arguing that the practical and ethical trade-offs inherent in complex decision-making processes are unavoidable. They propose a shift towards understanding fairness not as a static, absolute mathematical state to be achieved, but as a relational notion. This perspective advocates for evaluating the fairness of an algorithmic system relative to other viable options and explicitly considering the trade-offs between various, often competing, objectives. Applying this to 2011 Home Mortgage Disclosure Act (HMDA) data, they emphasize the limitations of standard mathematical metrics in capturing the nuanced ethical dilemmas faced by lenders at face value. This is particularly important for lending given biases embedded within legacy data, and the difficulty of untangling legitimate risk proxies from potentially discriminatory proxies for protected characteristics, e.g., race. The relational approach encourages an evaluation focused on achieving a model as close to equilibrium between desired outcomes, such as enhancing financial access, and potential negative impacts on specific groups. It marks a significant conceptual move, steering the literature away from the pursuit of a singular, perfect mathematical fairness definition towards a more pragmatic, context-dependent framework centered on navigating and managing competing values.
@@
Algorithmic fairness metrics are broadly categorized under two main paradigms: individual fairness and group fairness (anderson2025algorithmic). Individual fairness (IF) adheres to the principle that ”similar individuals should be treated similarly”. In this context, similarity is determined based on task-relevant characteristics, independent of protected attributes. For instance, two loan applicants with comparable financial profiles should ideally receive similar loan decisions. However, defining an appropriate, context-specific similarity metric remains a significant challenge, and subjective or data-driven definitions risk encoding existing biases. Group fairness (GF), conversely, focuses on achieving parity in outcomes or statistical measures across different demographic groups, typically defined by protected attributes such as race, sex, religion, national origin, or familial status (gohar2023survey). The objective is often to equalize specific rates, like approval rates or error rates, between these groups (zhou2022group).
@@
A critical tension exists between these two paradigms: satisfying group fairness criteria does not guarantee individual fairness, and enforcing group-level parity can sometimes necessitate treating similar individuals differently based on their group membership, directly conflicting with the IF principle (zhou2022group). This conflict often reflects a deeper ethical divergence between ensuring procedural consistency for individuals (treating like cases alike) and pursuing distributive justice to rectify group-level disparities, which may be rooted in historical bias reflected in the data. Furthermore, focusing solely on isolated protected attributes can obscure intersectional discrimination experienced by individuals with multiple identities (e.g., race and gender) (gohar2023survey).
@@
To address identified biases and strive towards fairness goals, various intervention techniques have been developed, typically categorized according to the stage of the machine learning pipeline at which they are applied: pre-processing, in-processing, and post-processing (dieng2025algorithmic).
  1. (1)
@@
    Pre-processing techniques: These techniques involve modifying the training data before a model is trained, removing or mitigating biases present in the data itself. Examples include suppressing sensitive attributes or their proxies, re-labeling historically biased outcomes, re-weighting data points to balance group representation, or learning on transformed data representations designed to be fair nielsen2020practical. The main advantage is that these techniques are model-agnostic and tackle bias at its source; some research suggests that successful pre-processing to approximate an unbiased world could resolve fairness-accuracy trade-offs (lein). However, pre-processing can significantly alter the original data, simple suppression is often ineffective due to correlations, and relabeling can be subjective and practically challenging.
  2. (2)
@@
    In-processing techniques: They integrate fairness considerations directly into the model training process, typically by modifying the learning algorithm or its objective function (lein). Common approaches include adding fairness-related regularization terms to the loss function to penalize unfair predictions (petersen2021post) or imposing fairness constraints during optimization. In-processing methods can potentially achieve a better balance between fairness and accuracy by optimizing them jointly (zemel2024group). These in-processing methods include in-training approaches such as adversarial debiasing (zhang_2018) and distribution matching (QuadriantoSharmanska2017) techniques. The latter of which was applied to mortgage underwriting and pricing models in a study which found that distribution matching reduced racial disparities in loan approvals by 13% and decreased borrowing costs for Black and Hispanic applicants by 20% without compromising system accuracy (merrill2024fairness). These methods promote fairness while maintaining efficiency, however, drawbacks include often being model-specific, requiring access to and modification of the training pipeline, and incurring the computational cost of retraining (zemel2024group).
  3. (3)
@@
    Post-processing techniques: They adjust the outputs of an already trained model without altering the model itself (zemel2024group). Examples include setting different classification thresholds for different groups or learning a separate transformation function for the model’s scores. The primary appeal lies in being model-agnostic, avoiding costly retraining, and applicability to deployed or black-box systems. However, their effectiveness is constrained by the information contained in the original model’s predictions, they may be less impactful than in-processing methods, and can sometimes improve group fairness at the expense of individual fairness or accuracy.
@@
The choice among these intervention strategies is not merely technical but involves navigating practical constraints—such as access to data, control over the model training process, and computational resources—alongside strategic goals regarding the desired fairness definition and tolerance for accuracy trade-offs (lein).
2.2. Quantifying Fairness: A Review of Mathematical Fairness Metrics
@@
In conjunction with theoretical fairness discussions, the field has created a toolkit of fairness metrics to quantify and describe variability between different subgroups within a dataset—-in the housing sector this usually includes demographic parity, standardized mean deviation, disparate/adverse impact, and mean absolute deviation. A common definition for fairness in this context points to opportunity access rather than to outcomes themselves lee21. Equalization of evaluation metrics attempt to formalize this belief that a model’s predictive accuracy should not vary between racial groups. The most utilized and cited of these include equal opportunity (EOP), false positive error rate balance, equal odds, positive predictive parity, positive class balance, and negative class balance.
@@
Equal opportunity, as discussed in Hardt2016 ensures equal chances of loan or housing approval for all credit-worthy applicants, focusing on minimizing false negative rates across demographics. Chouldechova2017’s work on false positive error rate balance seeks to ensure that similar non-credit worthy applicants are equally likely to be denied loans, preventing bias in loan disapproval. Equal odds requires algorithms to achieve equal true positive and false positive rates across groups, ensuring accuracy in identifying credit-worthy and non-worthy applicants, irrespective of demographic background Hardt2016. Positive predictive parity, aims for equal probabilities of correct loan repayment predictions across different groups, ensures that the algorithm’s trust is uniformly distributed, preventing skewed good faith predictions that favor specific demographics. Finally, Kleinberg2016’s work on positive class balance and negative class balance ensures that the confidence in predictions—whether for loan repayment or default—is fairly balanced across groups, reflecting an objective assessment of an applicant’s financial behavior.
@@
Collectively, fairness metrics describe data variance and facilitate the evaluation of outcome distributions. For developers, policymakers, and stakeholders navigating the concerns associated with algorithmic decision-making, they act as a key benchmarking strategy for documenting how ’fairness’ within an algorithm is distributed across subgroups in a dataset. However, as remarked in lee21, while adept at quantifying disparities, it is also essential to note that these traditional metrics solely reflect relative comparisons between groups, similar individuals, or group error rates, without necessarily accounting for the biases embedded within the data inherently. This can result then in different effect distributions if comparing the conclusions suggested by metrics with real-world consequences. This is demonstrated in noriega2020algorithmic’s applied case study in algorithmic targeting of social policy in Columbia and Costa Rica. While finding overwhelming positive predictive improvements by using AI-based systems compared to the status quo, their analysis underscores that the degree of these improvements were not evenly dispersed. Error reduction ranged between 5.74% to 22.57% in some cases, across various urban-rural, geographic, gender, and familial divisions, highlighting the importance of documenting and testing fairness outcomes across segmented subgroups in high-impact use cases. Fairness metrics, though helpful in highlighting disparities, often present challenges in implementation, particularly in cases when the proxies for outcomes are closely intertwined with those for race, meaning they are best used when heavily contextualized to the data’s use case.
@@
2.3. Critical Evaluation and Adaptation of lee21: Strengths and Limitations of the Relational Trade-off Methodology
@@
While informed by the work of lee21, the research presented herein adapts certain assumptions employed in their analysis, based on limitations identified within their methodology and the specific requirements of this study. The methodology proposed in Section 5 of the paper offers a distinct approach centered on relational trade-offs. It involves several key steps: first, operationalizing relevant real-world objectives (e.g., defining and quantifying financial inclusion and negative impact on minority borrowers); second, building various predictive algorithms and computing metrics related to these objectives; third, identifying potential proxies for protected characteristics within the model features; and finally, selecting an algorithm by visualizing and evaluating the trade-offs between the competing objectives.
@@
This approach possesses notable strengths. It promotes a holistic and contextual view of fairness, moving beyond abstract mathematical conditions to evaluate algorithms within their specific application context and considering multiple, potentially conflicting, objectives. A significant advantage is its transparency regarding ethical trade-offs; it makes the inherent conflicts between values (like profit versus fairness) explicit and quantifiable, compelling decision-makers to confront and prioritize these trade-offs, aligning with broader calls for tools to evaluate AI trustworthiness (petersen2023risky). This focus on measurable impacts and comparing concrete algorithmic alternatives lends the methodology practical relevance and actionability for stakeholders, in this case lenders, aiding them in selecting algorithms that align with their ethical commitments and level of risk tolerance. Furthermore, the methodology explicitly incorporates steps to identify and analyze potential proxies for protected attributes, fostering a more profound understanding of how bias might operate, even when sensitive attributes are formally excluded.
@@
The methodology, however, also presents limitations. Its outcomes are highly dependent on the initial assumptions made when operationalizing objectives; how concepts like ”financial inclusion” are defined and measured significantly shapes the resulting analysis. This leads to potential subjectivity, as the interpretation of the trade-offs and the final algorithm choice rely heavily on the decision-maker’s values, potentially lacking the perceived objectivity of purely mathematical metrics. The analysis is also heavily dependent on data quality and availability; lee21 themselves noted limitations in the public HMDA data used in their case study, which lacked crucial variables like default outcomes and credit scores, impacting the robustness of the final trade-off evaluation. Moreover, quantifying all relevant factors (e.g., regulatory compliance, long-term societal effects, model interpretability) can be challenging, and the original study explored a limited scope of algorithms and stakeholder perspectives. Ultimately, the framework functions less as a definitive solution to fairness problems and more as a structured process for deliberation. It requires stakeholders to make explicit value judgments at multiple stages—defining objectives, assessing proxies, prioritizing trade-offs—thereby facilitating, rather than automating, ethical decision-making. Its strength lies in clarifying these choices, while its weakness is its reliance on the quality of the inputs and the deliberation itself.